OpenStudy (samsan9):

find f(2) and f(a+h) when f(x)=3x^2+2x+4

OpenStudy (anonymous):

just substitute 2 instead of x and a + h instead of x

OpenStudy (anonymous):

Plug 2 into the equation and that will give you the value of f(x). Do the same with(a+h)

OpenStudy (samsan9):

i got 20 for the first one but you lost me at the (a+h)

OpenStudy (anonymous):

Where ever there is an x in the equation replace it with (a+h).

OpenStudy (anonymous):

3 ( a+h)^2 +2 (a+h)+4 then get the brackets out of here :d

OpenStudy (anonymous):

f(?) the (?) simply telles you what value to input in the equation.

OpenStudy (samsan9):

so is it f(2)=20 and f(a+h)= 3(a+h)^2+2(a+h)+4?

OpenStudy (aum):

f(2) = 20 is correct. But you have to simplify 3(a+h)^2+2(a+h)+4 = 3(a^2 + 2ah + h^2) + 2a + 2h + 4 = 3a^2 + 6ah + 3h^2 + 2a + 2h + 4

OpenStudy (samsan9):

so that is the correct answer wow that was long but thank you :D

OpenStudy (aum):

You are welcome.